Records of a health insurance company show that 30% of its policyholders over 50 submit a claim during the year. 15 policyholders over 50 are selected at random; what is the probability that at least 10 will submit a claim during the year? 

What is the probability that 4 will submit a claim during the year?

 How many do you expect to submit a claim?

What is the standard deviation?



1
Expert's answer
2021-10-19T13:16:09-0400

This is a binomial distribution with n=15, p=0.3.

"P(X\ge10)=\Sigma_{i=10}^{15}P(X=i)=\Sigma_{i=10}^{15}C_{15}^i0.3^i(1-0.3)^{15-i}=0.0037."


"P(X=4)=C_{15}^40.3^4(1-0.3)^{11}=0.2186."


"E(X)=np=15*0.3=4.5."


"\sigma=\sqrt{np(1-p)}=\sqrt{15*0.3*0.7}=\sqrt{3.15}=1.7748."